OpenStudy (kinggeorge):

[SOLVED] While I'm not terribly busy, have this as a challenge problem. What are the next two numbers in this sequence of four digit positive integers? \[8741, 7632, 6552, 9963, 6642, \_\_\_\_, \_\_\_\_\]

OpenStudy (shubhamsrg):

i have heard about this before ans is 7641,7641 please confirm

OpenStudy (shubhamsrg):

i also know the deep theory behind it,though roughly only..

OpenStudy (anonymous):

@shubhamsrg I would like to know the reason

OpenStudy (shubhamsrg):

i'd want to confirm @KingGeorge to confirm the ans first.. otherwise am 99% convinced am correct.. i'll message you the reason..

OpenStudy (anonymous):

Looks like u r right @shubhamsrg

OpenStudy (shubhamsrg):

hmm.. :)

OpenStudy (kinggeorge):

We have a winner! Excellent work. I was wondering how long it would be until someone noticed the pattern :P

OpenStudy (shubhamsrg):

thank you :)

OpenStudy (kinggeorge):

For those wondering what the pattern is, you take the starting number, order the digits in increasing order to create one new number, and order the digits in decreasing order to create a second number. Then, you subtract the smaller number from the larger number, and repeat the process with your new number. As I did above, 8741-1478=7263 7632-2367=5265 6552-2556=3996 9963-3699=6264 6642-2466=4176 7641-1467=6174 Curiously enough, if you do this pattern for any positive 4-digit number, you always get 6174 after 7 steps or less. See here for more info. http://en.wikipedia.org/wiki/6174_(number)

Parth (parthkohli):

Wow, that's a nice conjecture! :) Can it be prove? @KingGeorge

OpenStudy (kinggeorge):

@ParthKohli I do not have the required knowledge to prove it other than through brute force. However, it can be proven with brute force. I should also qualify that there are some numbers (1111, 1211, 4443, 4444, etc) that do not reach 6174, and instead reach 0.

Parth (parthkohli):

Proved* Thanks :-)